OpenStudy (anonymous):

One angle of 180° is divided in to angles. The ratio of the angles is 4/5. Find the greatest angle?

OpenStudy (anonymous):

that is what I get, however, the possible answers for the greatest angle are: 100,90,80,50

OpenStudy (anonymous):

Because the ratio of the angles is 4 to 5, the math statement is going to be: \[4x+5x = 180\]

OpenStudy (anonymous):

x is going to be 20, so 4x = 80, and 5x = 100

OpenStudy (anonymous):

4x+5x=180 9x=180 x=20 4(20)+5(20)=180 80+100=180 little angle 80, big angle 100

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome!